# A Minimal Model for Emergent Collective Behaviors in Autonomous Robotic Multi-Agent Systems

### v1

@@ 4 The Cognitive Adaptation Mechanism
Cognitive models offer a powerful paradigm for the development of autonomous robotic collectives capable of context-aware and adaptive decision-making in dynamic environments. Collective behaviors in such systems emerge from the context-sensitive modulation of individual actions, driven by internal objectives aligned with adaptive goals, such as survival, task execution, or resource optimization.
We frame the transition between flocking and swarming as an emergent cognitive process wherein agents continuously re-evaluate and prioritize competing behavioral imperatives, namely, the maintenance of group cohesion versus the pursuit of environmental exploration. This formulation draws inspiration from natural systems, where animal groups exhibit collective behavioral shifts between flocking during migration for energy efficiency and dispersal during foraging for resource acquisition.
@@ 5.3 Asymptotic stability
Then, all agents asymptotically achieve their desired relative positions and velocities defined by the offset vectors $\mathbf{p}_{ij}$ and $\mathbf{v}_{ij}$.
@@ 6.1 Minimal model
First, we validate the model by showing that it generates various motion patterns under different settings, as illustrated in Fig. 1. Emerging collective behaviors are flocking (Fig. 1(1(a))), vortexing (Fig. 1(1(b))), and swarming (Fig. 1(1(c))).
@@
and present the results in Fig. 2. Figs. 2(2(a))–(2(d)) correspond to the flocking scenario, showing the evolution of position and velocity errors along the $x$- and $y$-axes. The plots exhibit rapid convergence of position errors, indicative of cohesive motion. Velocity convergence, however, remains partial due to the nonzero kinetic offset $\eta_{i}=3$, which permits controlled velocity divergence among neighboring agents. Figs. 2(2(e))–(2(h)) represent the vortexing scenario. In this case, both position and velocity errors rapidly diminish once the vortex pattern stabilizes, confirming the emergence of coordinated rotational motion. Finally, Figs. 2(2(i))–(2(l)) depict the swarming behavior. While the position errors converge swiftly—ensuring group cohesion—the velocity errors remain relatively large, consistent with the high kinetic offset $\eta_{i}=12$, which facilitates significant velocity heterogeneity and promotes exploratory dynamics.
@@
The typical motion patterns and behavioral phase transition diagram shown in Fig. 3(3(a)), the corresponding radii of aggregation in Fig. 3(3(b)), and the average speeds in Fig. 3(3(c)) provide an overview of our model for various population sizes $n$. In particular, as $n$ increases, the behavioral phase transition becomes smoother and eventually converges to random motion directions at the individual level. In summary, Fig. 3 demonstrates that $\eta_{i}$ induces the behavioral phase transition in collective motion, leading to the emergence of flocking, vortexing, and swarming behaviors.
@@
Vortexing behavior does not emerge explicitly through the kinetic offset parameter tuning across arbitrary population sizes. However, as illustrated in Fig. 3(3(a)), it consistently arises as a typical collective pattern in small-scale populations ($n<10$). Future work may explore model extensions or variations that allow vortexing behavior on broader population scales.
@@
To suppress the emergence of structured disorder—manifested as vortex-like motion, as illustrated in Fig. 1(1(b))—and instead promote disordered yet cohesive swarming behavior in low-density settings (e.g. $n<10$), the parameter $\beta$ can be effectively increased. A larger $\beta$ amplifies the alignment weight $\phi(\cdot)$, thereby preventing the formation of structured motion patterns and facilitating the globally disordered dynamics characteristic of swarm-like behavior. For example, increasing $\beta$ from $1$ to $5$ transitions the collective behavior from that shown in Fig. 1(1(b)) to the swarm-like behavior depicted in Fig. 1(1(c)).
@@
Figure 4: Phase transition diagram illustrating the alignment score $h$ as a function of the kinetic offset $\eta_{i}$ and population size $n$. Contour levels represent the degree of alignment achieved by the collective for varying combinations of $\eta_{i}$ and $n$, revealing critical transition regions where collective behavior shifts from disordered to ordered states. The uncolored region corresponds to the critical $\eta_{i}$ values, beyond which aggregation in a population of size $n$ breaks down, as reflected in the dispersion radii shown in Fig. 3(3(b)).
@@
In Fig. 5, we plot the minimum distance, average distance, and radii of aggregation for the collective motion of $n=100$ agents evolving over $t=30$ s, with gradually increasing values of $\delta_{i}$. Figs. 5(5(a))–5(5(c)) correspond to $\eta_{i}=3$, which produces a relatively ordered collective motion, exhibiting flocking behavior. As seen in Fig. 5(5(a)), a nonzero $\delta_{i}$ prevents inter-agent collisions and results in a relatively stable separation distance. Both the average distances and radii of aggregation increase as $\delta_{i}$ increases, with $\delta_{i}=0.5$ and $1$ leading to the smoothest flocking behavior. Figs. 5(5(d))–5(5(f)) correspond to $\eta_{i}=21$, which leads to disordered collective motion, exhibiting swarming behavior. Although the evolution of disorder engenders random motion directions at the individual level and consequently results in difficulties in establishing a stable minimum distance, collision avoidance is nevertheless preserved, as observed from Fig. 5(5(d)). Notably, $\delta_{i}=0.5$ and $1$ promote safer and more stable swarming behavior, as the fluctuations in minimum distances are minimized. Accordingly, Fig. 5 illustrates that the degree of aggregation in collective motion is directly modulated by the agents’ propensity to aggregate, as parameterized by the spatial offset $\delta_{i}$. Moreover, a nonzero $\delta_{i}$ effectively prevents collisions, with the desired results observed for $\delta_{i}=0.5$ and $1$ in both flocking and swarming behaviors.
@@ 6.2 Energy-driven adaptation
To evaluate the cognitive model with the energy-driven adaptation mechanism, we define the parameters $\delta_{\text{min}}$, $\delta_{\text{max}}$, $\eta_{\text{min}}$, and $\eta_{\text{max}}$, allowing agents to adjust their policies $\delta_{i}$ and $\eta_{i}$ based on their energy level $E_{i}$. From Fig. 3(3(b)), we define $\eta_{\max}=13$ for $n=2$, $\eta_{\max}=26$ for $n=3$, and $\eta_{\max}=33$ for $n>3$. However, to ensure robustness in the presence of varying $\eta_{i}$ among neighbors, a more conservative choice is $\eta_{\max}=15$, as extreme values may not sustain aggregation. Additionally, Fig. 5 provides an intuitive basis for determining $\delta_{\max}$. Specifically, $\delta_{\max}=2.5$ for flocking and $\delta_{\max}=1.5$ for swarming. Based on these observations, we set $\delta_{\max}=2$ and $\delta_{\min}=0.5$ as reasonable values, ensuring that agents remain within their interaction zone radius $r_{i}$.
We run the model with $n=20$ agents, both with and without the adaptation mechanism enabled. Fig. 6(6(a))–(6(c)) presents the time histories of $\eta_{i}$, $\delta_{i}$, and $E_{i}$ for flocking behavior in a compact formation, where agents exhibit strong alignment and maintain close proximity to their neighbors. Fig. 6(6(e))-(6(g)) displays the corresponding time histories for swarming behavior, where agents are driven to explore the environment. Finally, Fig. 6(6(i))-(6(k)) illustrates the time histories of $\eta_{i}$, $\delta_{i}$, and $E_{i}$ when agents dynamically adapt $\eta_{i}$ and $\delta_{i}$ based on their energy levels $E_{i}$. The snapshots of collective motion under the energy-driven adaptive mechanism are depicted in Fig. 7, illustrating how agents individually adjust their $\delta_{i}$ and $\eta_{i}$ policies in response to their own and their neighbors’ energy levels, thereby shaping the dynamics of collective behavior.
From Fig. 6(6(c)), Fig. 6(6(g)), and Fig. 6(6(k)), three key observations emerge: (1) flocking in a compact formation conserves energy, (2) swarming behavior incurs higher energy expenditure, and (3) the energy-driven adaptation mechanism effectively regulates energy efficiency by transitioning agents from energy-intensive swarming to energy-conserving flocking when their energy drops below a threshold. This highlights the mechanism’s ability to optimize agent behavior under energy constraints.
@@
Figure 7: Snapshots of collective motion under the energy-driven adaptive mechanism. Agents dynamically adjust their $\delta_{i}$ and $\eta_{i}$ policies in response to their own and neighbors’ energy levels, shaping the emergent collective behavior. Initialized with randomized positions and velocities in (a), the collective behavior transitions to swarming in (b) to (d) and evolves into flocking in (e) to (h).

### v2

@@ 5 The Cognitive Adaptation Mechanism
Cognitive models enable autonomous robotic collectives to adaptively respond to dynamic environments through context-sensitive actions aligned with goals like survival or resource optimization. We frame the transition between flocking and swarming as an emergent cognitive process wherein agents continuously re-evaluate and prioritize competing behavioral imperatives, namely, the maintenance of group cohesion versus the pursuit of environmental exploration. This formulation draws inspiration from natural systems, where animal groups exhibit collective behavioral shifts between flocking during migration for energy efficiency and dispersal during foraging for resource acquisition.
@@ 6.3 Asymptotic stability
Then, all agents asymptotically achieve their desired relative positions and velocities defined by the offset vectors $\mathbf{p}_{ij}$ and $\mathbf{v}_{ij}$. The analysis assumes a fixed topology for tractability, while the model inherently features dynamic neighbor selection, resulting in a time-varying topology. This is addressed by applying the fixed-topology stability result at each time step, treating the current state as a new initial condition, provided connectivity is preserved. Hence, convergence to the desired relative positions and velocities remains valid.
@@ 7.1 Minimal model
First, we validate the model by showing that it generates various motion patterns under different settings, as illustrated in Fig. 3. Emerging collective behaviors are flocking (Fig. 3(3(a))), vortexing (Fig. 3(3(b))), and swarming (Fig. 3(3(c))).
@@
and present the results in Fig. 4. Figs. 4(4(a))–(4(d)) correspond to the flocking scenario, showing the evolution of position and velocity errors along the $x$- and $y$-axes. The plots exhibit rapid convergence of position errors, indicative of cohesive motion. Velocity convergence, however, remains partial due to the nonzero kinetic offset $\eta_{i}=3$, which permits controlled velocity divergence among neighboring agents. Figs. 4(4(e))–(4(h)) represent the vortexing scenario. In this case, both position and velocity errors rapidly diminish once the vortex pattern stabilizes, confirming the emergence of coordinated rotational motion. Finally, Figs. 4(4(i))–(4(l)) depict the swarming behavior. While the position errors converge swiftly—ensuring group cohesion—the velocity errors remain relatively large, consistent with the high kinetic offset $\eta_{i}=12$, which facilitates significant velocity heterogeneity and promotes exploratory dynamics.
@@
The typical motion patterns and behavioral phase transition diagram shown in Fig. 5(5(a)), the corresponding radii of aggregation in Fig. 5(5(b)), and the average speeds in Fig. 5(5(c)) provide an overview of our model for various population sizes $n$. In particular, as $n$ increases, the behavioral phase transition becomes smoother and eventually converges to random motion directions at the individual level. In summary, Fig. 5 demonstrates that $\eta_{i}$ induces the behavioral phase transition in collective motion, leading to the emergence of flocking, vortexing, and swarming behaviors.
@@
Vortexing behavior does not emerge explicitly through the kinetic offset parameter tuning across arbitrary population sizes. However, as illustrated in Fig. 5(5(a)), it consistently arises as a typical collective pattern in small-scale populations ($n<10$). Future work may explore model extensions or variations that allow vortexing behavior on broader population scales.
@@
To suppress the emergence of structured disorder—manifested as vortex-like motion, as illustrated in Fig. 3(3(b))—and instead promote disordered yet cohesive swarming behavior in low-density settings (e.g. $n<10$), the parameter $\beta$ can be effectively increased. A larger $\beta$ amplifies the alignment weight $\phi(\cdot)$, thereby preventing the formation of structured motion patterns and facilitating the globally disordered dynamics characteristic of swarm-like behavior. For example, increasing $\beta$ from $1$ to $5$ transitions the collective behavior from that shown in Fig. 3(3(b)) to the swarm-like behavior depicted in Fig. 3(3(c)).
@@
Figure 6: Phase transition diagram illustrating the alignment score $h$ as a function of the kinetic offset $\eta_{i}$ and population size $n$. Contour levels represent the degree of alignment achieved by the collective for varying combinations of $\eta_{i}$ and $n$, revealing critical transition regions where collective behavior shifts from disordered to ordered states. The uncolored region corresponds to the critical $\eta_{i}$ values, beyond which aggregation in a population of size $n$ breaks down, as reflected in the dispersion radii shown in Fig. 5(5(b)).
@@
In Fig. 7, we plot the minimum distance, average distance, and radii of aggregation for the collective motion of $n=100$ agents evolving over $t=30$ s, with gradually increasing values of $\delta_{i}$. Figs. 7(7(a))–7(7(c)) correspond to $\eta_{i}=3$, which produces a relatively ordered collective motion, exhibiting flocking behavior. As seen in Fig. 7(7(a)), a nonzero $\delta_{i}$ prevents inter-agent collisions and results in a relatively stable separation distance. Both the average distances and radii of aggregation increase as $\delta_{i}$ increases, with $\delta_{i}=0.5$ and $1$ leading to the smoothest flocking behavior. Figs. 7(7(d))–7(7(f)) correspond to $\eta_{i}=21$, which leads to disordered collective motion, exhibiting swarming behavior. Although the evolution of disorder engenders random motion directions at the individual level and consequently results in difficulties in establishing a stable minimum distance, collision avoidance is nevertheless preserved, as observed from Fig. 7(7(d)). Notably, $\delta_{i}=0.5$ and $1$ promote safer and more stable swarming behavior, as the fluctuations in minimum distances are minimized. Accordingly, Fig. 7 illustrates that the degree of aggregation in collective motion is directly modulated by the agents’ propensity to aggregate, as parameterized by the spatial offset $\delta_{i}$. Moreover, a nonzero $\delta_{i}$ effectively prevents collisions, with the desired results observed for $\delta_{i}=0.5$ and $1$ in both flocking and swarming behaviors.
@@
7.2 Cluttered environment
A simulation of 10 agents in a $100\times 100\,\text{m}$ environment with obstacles was conducted to validate the target-directed collective behavior model in a cluttered environment. The agents aimed to reach a target at $(90,90)\,\text{m}$ while avoiding three circular obstacles located at $(25,30)\,\text{m}$, $(50,40)\,\text{m}$, and $(90,80)\,\text{m}$, each with a radius of $5\,\text{m}$ , as shown in Fig. 8. The parameters were set as follows: $r_{i}=10\,\text{m}$, $\delta_{i}=1$, $\eta_{i}=0.5$, $\alpha=2$, $\beta=1$, $\kappa_{i}=0.5$, $c_{i}=15\,\text{m}$, and $\sigma=3$. Over a $40\,\text{s}$ simulation, the agents demonstrated cohesive, collision-free motion, effectively balancing aggregation, alignment, and obstacle avoidance. The flock temporarily fragmented near the second obstacle but later regrouped and successfully reached the target. The choice of $\eta_{i}=0.5$ accounts for environmental noise, ensuring stable alignment under realistic conditions.
Figure 8: Snapshots of 10 agents navigating a $100\times 100\,\text{m}$ environment toward a target at $(90,90)\,\text{m}$, avoiding obstacles at $(25,30)\,\text{m}$, $(50,40)\,\text{m}$, and $(90,80)\,\text{m}$, each with a $5\,\text{m}$ radius. With initial positions in $0,10]\,\text{m}$, subfigures depict: (a) encounter with the obstacle at $(25,30)\,\text{m}$, (b) encounter with the obstacle at $(50,40)\,\text{m}$, (c) encounter with the obstacle at $(90,80)\,\text{m}$, and (d) arrival at the target. Trajectories are shown in green, agent positions as red circles, velocities as red arrows, the target as a blue star, and obstacles as black circles.
@@ 7.3 Energy-driven adaptation
To evaluate the cognitive model with the energy-driven adaptation mechanism, we define the parameters $\delta_{\text{min}}$, $\delta_{\text{max}}$, $\eta_{\text{min}}$, and $\eta_{\text{max}}$, allowing agents to adjust their policies $\delta_{i}$ and $\eta_{i}$ based on their energy level $E_{i}$. From Fig. 5(5(b)), we define $\eta_{\max}=13$ for $n=2$, $\eta_{\max}=26$ for $n=3$, and $\eta_{\max}=33$ for $n>3$. However, to ensure robustness in the presence of varying $\eta_{i}$ among neighbors, a more conservative choice is $\eta_{\max}=15$, as extreme values may not sustain aggregation. Additionally, Fig. 7 provides an intuitive basis for determining $\delta_{\max}$. Specifically, $\delta_{\max}=2.5$ for flocking and $\delta_{\max}=1.5$ for swarming. Based on these observations, we set $\delta_{\max}=2$ and $\delta_{\min}=0.5$ as reasonable values, ensuring that agents remain within their interaction zone radius $r_{i}$.
We run the model with $n=20$ agents, both with and without the adaptation mechanism enabled. Fig. 9(9(a))–(9(c)) presents the time histories of $\eta_{i}$, $\delta_{i}$, and $E_{i}$ for flocking behavior in a compact formation, where agents exhibit strong alignment and maintain close proximity to their neighbors. Fig. 9(9(e))-(9(g)) displays the corresponding time histories for swarming behavior, where agents are driven to explore the environment. Finally, Fig. 9(9(i))-(9(k)) illustrates the time histories of $\eta_{i}$, $\delta_{i}$, and $E_{i}$ when agents dynamically adapt $\eta_{i}$ and $\delta_{i}$ based on their energy levels $E_{i}$. The snapshots of collective motion under the energy-driven adaptive mechanism are depicted in Fig. 10, illustrating how agents individually adjust their $\delta_{i}$ and $\eta_{i}$ policies in response to their own and their neighbors’ energy levels, thereby shaping the dynamics of collective behavior.
From Fig. 9(9(c)), Fig. 9(9(g)), and Fig. 9(9(k)), three key observations emerge: (1) flocking in a compact formation conserves energy, (2) swarming behavior incurs higher energy expenditure, and (3) the energy-driven adaptation mechanism effectively regulates energy efficiency by transitioning agents from energy-intensive swarming to energy-conserving flocking when their energy drops below a threshold. This highlights the mechanism’s ability to optimize agent behavior under energy constraints.
@@
Figure 10: Snapshots of collective motion under the energy-driven adaptive mechanism. Agents dynamically adjust their $\delta_{i}$ and $\eta_{i}$ policies in response to their own and neighbors’ energy levels, shaping the emergent collective behavior. Initialized with randomized positions and velocities in (a), the collective behavior transitions to swarming in (b) and evolves into flocking in (c) to (d).